A courier service company has found that their delivery time of parcels to clients is approximately normally distributed with a mean delivery time of 30 minutes and a variance of 25 minutes (squared). Required: a) What is the probability that a randomly selected parcel will take more than 26 minutes to deliver? (2) b) What is the minimum delivery time (minutes) for the 2.5% of parcels with the longest time to deliver? (4


1
Expert's answer
2021-05-31T10:39:56-0400

Let "X=" delivery time of parcels: "X\sim N(\mu, \sigma^2)."

Given "\mu=30\ min, \sigma^2=25\ min^2"

a)


"P(X>26)=1-P(X\leq 26)"

"=1-P(Z\leq\dfrac{26-30}{5})=1-P(Z\leq-0.8)"

"\approx0.788147"

b)


"P(X<x)=P(Z<\dfrac{x-30}{5})=0.025"

"\dfrac{x-30}{5}\approx-1.96"

"x\approx 20.2\ min"